OpenStudy (anonymous):

two charges of -6.00 micro coulombs and +8.00 micro coulombs attract each other with a force of 3.0x10^3N in a vacuum. Calculate the distance between the particles?

OpenStudy (anonymous):

Force= k*q1.q2 / r2 k =8.9 so calculate

OpenStudy (anonymous):

q1 and q2 are the charges

OpenStudy (anonymous):

f=kq1\times q2divr^2\],so yr answer is6.9*10^3

OpenStudy (anonymous):

srry yr answer is 6.9 metre

OpenStudy (anonymous):

Using the Coulomb's Law F=K*q1*q1/r^2 where F is the electrical force between the particles, q,q1 the particles' charge, r the distance between the particles and K=9*10^9 N*m^2/C^2 we solve the equation and we find r is about 8.45m.

OpenStudy (anonymous):

My bad.. we find r^2=0.00072m so r=0.027m (I accidentally screwed the calculations and found r^2=72m so that's how i got the wrong calculation).

OpenStudy (anonymous):

k=9.8,i have used